OpenStudy (anonymous):

how do i write an equation of a circle that has a diameter with the given end points. (-5,3) and (3,3)

OpenStudy (anonymous):

If you take the diameter and use the midpoint equation, you will find the center of the circle. Then you can calculate the radius by taking half of the diameter length. Then you can fill in the parts of the circle equation :)

OpenStudy (anonymous):

mid point of the diameter: (-5-3/2,3-3/2) =(-4,0) (x+4) 2 +y 2 =58

OpenStudy (anonymous):

^ idk if im correct but i got (x+1)^2 + (y-3)^2= 16